OpenStudy (anonymous):

How do I find the directrix and focal length of a parabola? For example, y=1/12x^2

OpenStudy (stamp):

any point on a parabola is equidistant to the focus and the directrix

OpenStudy (anonymous):

But how do I find the directrix...?

OpenStudy (stamp):

there is probably a formula

OpenStudy (anonymous):

Ok, thanks.

OpenStudy (amistre64):

i believe the form is properly: 4py =x^2

OpenStudy (anonymous):

Oh, it is D:?

OpenStudy (amistre64):

y = 4x^2 has a focal point at (0, 1/16) if we assume y=4px^2, that would make the focal point at (0,1) ...

OpenStudy (anonymous):

Hm...that does seem to make sense. I put an example in my question. How would that work out?

OpenStudy (amistre64):

y=1/12x^2 ; *12 12y = x^2 let 4p = 12 , p = 3 gives the distance we need from the vertex to find the focus and the directrix

OpenStudy (amistre64):